# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ Abstract
We present new concentration inequalities for either martingale dependent or exchangeable random symmetric matrices under a variety of tail conditions, encompassing now-standard Chernoff bounds to self-normalized heavy-tailed settings. These inequalities are often randomized in a way that renders them strictly tighter than existing deterministic results in the literature, are typically expressed in the Loewner order, and are sometimes valid at arbitrary data-dependent stopping times. Along the way, we explore the theory of positive semidefinite supermartingales and maximal inequalities, a natural matrix analog of scalar nonnegative supermartingales that is potentially of independent interest.
@@ 1 Introduction
Bounding the tail probability of the sum of independent and identically distributed (i.i.d.) real-valued random variables is one of the most crucial tasks in the nonasymptotic theory of statistics, for which a copious collection of methods has flourished [Boucheron et al., 2013]. This paper concerns concentration inequalities for sums of random symmetric matrices under martingale dependence or exchangeability (both generalizing the i.i.d. assumption).
Notable previous results include the study of Rademacher and random graph adjacency matrices by Oliveira [2009, 2010]; those in a series of seminal papers by Tropp [2012, 2015], who famously ports the techniques of cumulant generating functions (CGFs) and Chernoff-type bounds, which are central in scalar concentration, into the matrix regime; some new bounds by Mackey et al. [2014] using exchangeability; and a recent study on “anytime-valid” concentration inequalities by Howard et al. [2020, 2021] which improves and unifies several matrix Chernoff-type bounds. We return to several of these works later in the current paper.
To set the stage for our exposition, we quote a celebrated result by Mackey et al. [2014, Corollary 4.2] (and also Tropp [2012, Theorem 1.3]) which states that, if $d\times d$ random i.i.d. symmetric matrices $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ satisfy $(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})^{2}\preceq\mathbf{B}$ (meaning that $\mathbf{B}-(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})^{2}$ is positive semidefinite) almost surely, then
$\mathbb{P}\left(\lambda_{\max}\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}-\mathbb{E}\mathbf{X}_{1}\right)\geqslant\sqrt{\frac{2\lambda_{\max}(\mathbf{B})\log(d/\alpha)}{n}}\right)\leqslant\alpha.$ (MHI)
Above, $\lambda_{\max}(\,\cdot\,)$ refers to the largest eigenvalue of a symmetric matrix, a typical scalarization of matrix-valued tails. When $d=1$, it recovers, with the same optimal constant, the scalar Hoeffding’s inequality as a special case; hence its name “matrix Hoeffding’s inequality”.
In this paper, we present numerous new techniques to strengthen this type of matrix concentration inequalities. Each of our new bounds enjoys a combination of some of the following virtues:
@@
    Anytime validity. We derive new inequalities that hold not only for a fixed number ($n$) of matrices but also a stopping time number of them [Howard et al., 2020].
@@
    Randomization. We losslessly tighten many of the previously known inequalities by randomly decreasing the tail threshold (decreasing the $a$ in (MHI) on the left hand side but not the right hand side), extending a technique that was recently explored by Ramdas and Manole [2023] for scalars.
@@
    Exchangeability. We go beyond the i.i.d. or martingale difference assumptions on random matrices and derive bounds that allow them to just be exchangeable instead.
@@
    Tighter (or no) scalarization. In some bounds we use the trace functional $\mathrm{tr}(\,\cdot\,)$ which is larger than the usual $\lambda_{\max}(\,\cdot\,)$ (e.g. in (MHI)), and introduce some bounds completely in the matrix order $\preceq$ without any scalarization, which is partly achieved by studying the maximal inequalities of matrix-valued sub- and super-martingales.
@@
    Heavy tails. Instead of assuming the matrices to be bounded or (sub)Gaussian, we sometimes allow them to only have a finite variance or $p$th moment, or be symmetric.
To showcase these virtues, let us pick two of the bounds we shall prove in the exposition that follows. Under the assumptions of the (MHI) above, our Theorem 5.5 states that
$\mathbb{P}\left(\lambda_{\max}\left(\frac{\gamma_{1}\mathbf{X}_{1}+\dots+\gamma_{\tau}\mathbf{X}_{\tau}}{\gamma_{1}+\dots+\gamma_{\tau}}-\mathbb{E}\mathbf{X}_{1}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{2}\lambda_{\max}(\mathbf{B})\sum_{i=1}^{\tau}\gamma_{i}^{2}}{\gamma_{1}+\dots+\gamma_{\tau}}\right)\leqslant\alpha$ (1)
for any predictable positive tuning parameters $\{\gamma_{n}\}$, stopping time $\tau$, and an independent random variable $U\sim\text{Unif}_{(0,1)}$. The above equips (MHI) with virtues 1 and 2 without compromises, since replacing $U$ with 1 and taking $\tau=n$, $\gamma_{1}=\dots=\gamma_{n}=\sqrt{\frac{2\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B})}}$ in above, one obtains the exact deviation bound $\sqrt{\frac{2\log(d/\alpha)\lambda_{\max}(\mathbf{B})}{n}}$ as in (MHI). A bound of very different flavor is that of our Theorem 4.17, which states that for exchangeable matrices $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ with variance $\mathbf{V}$,
$\mathbb{P}\left(\exists n\geqslant 1:\operatorname{abs}\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}-\mathbb{E}\mathbf{X}_{1}\right)\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ (2)
for any positive definite $\mathbf{A}$. This bound enjoys virtues 3, 4, and 5, sequentializing the matrix Chebyshev inequality $\mathbb{P}\left(\operatorname{abs}\left(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1}\right)\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ by Ahlswede and Winter [2002, Appendix C] to an infinite time horizon.
Outline.
The rest of the paper is organized as follows. In Section 2, we briefly review the technical background (and challenges) of matrix concentration bounds as well as a collection of classical and recent scalar bounds that anticipate our matrix extensions. We present in Section 3 our first kind of bounds, derived via randomized matrix Markov, Chebyshev, and Chernoff inequalities. We develop and apply the theory of matrix-valued supermartingales and backward submartingales in Section 4, further obtaining bounds for i.i.d. and exchangeable matrices. Finally, in Section 5, we revisit scalar supermartingales (and the scalar processes they dominate) constructed via matrices which yield some other randomized bounds as well, concluding with a discussion on the respective advantages of matrix- and scalar- based sequential tests in both theoretical and empirical dimensions.
2 Preliminaries
@@ 2.1 Matrix Analysis
Let $\mathcal{S}_{d}$ denote the set of all $d\times d$ real-valued symmetric matrices, which is the only class of matrices considered in this paper. [^1] We only briefly discuss the case with non-symmetric and non-square matrices in Appendix D. These matrices are denoted by bold upper-case letters $\mathbf{A},\mathbf{B}$, etc; whereas $d$ dimensional (column) vectors by bold lower-case letters $\mathbf{u},\mathbf{v}$, etc. For $I\subseteq\mathbb{R}$, we denote by $\mathcal{S}_{d}^{I}$ the set of all real symmetric matrices whose eigenvalues are all in $I$. $\mathcal{S}_{d}^{[0,\infty)}$, the set of positive semidefinite and $\mathcal{S}_{d}^{(0,\infty)}$, the set of positive definite matrices are simply denoted by $\mathcal{S}_{d}^{+}$ and $\mathcal{S}_{d}^{++}$ respectively. The Loewner partial order is denoted $\preceq$, where $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite, and $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite.
Any function $f:I\to J$ (where $I,J\subseteq\mathbb{R}$) is identified canonically with a function $f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$ in the following way: for a diagonal matrix $\mathbf{\Sigma}\in\mathcal{S}_{d}^{I}$, $f(\mathbf{\Sigma})$ is obtained by applying $f$ entrywise; otherwise it is obtained by applying $f$ after orthogonal diagonalization, $f(\mathbf{U}^{\mathsf{T}}\mathbf{\Sigma}\mathbf{U})=\mathbf{U}^{\mathsf{T}}f(\mathbf{\Sigma})\mathbf{U}$. Notable matrix-to-matrix operators that recur in this paper defined this way include the matrix exponential $\exp:\mathcal{S}_{d}\to\mathcal{S}^{++}_{d}$, the matrix logarithm $\log:\mathcal{S}^{++}_{d}\to\mathcal{S}_{d}$, the matrix absolute value [^2] We use $\operatorname{abs}(\,\cdot\,)$ instead of ||$\cdot$|| ⋅ | to distinguish it from the determinant and other norms. $\operatorname{abs}:\mathcal{S}_{d}\to\mathcal{S}^{+}_{d}$, and powers $\mathbf{X}\mapsto\mathbf{X}^{k}$ for any $k\in\mathbb{R}$. Take heed that usually $\exp(\mathbf{A}+\mathbf{B})\neq\exp\mathbf{A}+\exp\mathbf{B}$ and $\log\mathbf{A}\mathbf{B}\neq\log\mathbf{A}+\log\mathbf{B}$, but a trace equality holds on the latter:
Fact 1 (Trace Logarithm Identity).
$\mathrm{tr}\log\mathbf{A}\mathbf{B}=\mathrm{tr}\log\mathbf{A}+\mathrm{tr}\log\mathbf{B}$.
@@
It is rare to find matrix-to-matrix functions that are both monotone (i.e. preserving the $\preceq$ order) and convex (i.e. $f$ such that $f(\sum\lambda_{i}\mathbf{X}_{i})\preceq\sum{\lambda_{i}}f(\mathbf{X}_{i})$ if $\sum\lambda_{i}=1$ and each $\lambda_{i}\geqslant 0$; concavity is defined similarly). Scalar-to-scalar monotonicity ($I\to J$) does not imply matrix-to-matrix monotonicity ($\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$), and the same holds for convexity. For example, the matrix square $\mathbf{X}\mapsto\mathbf{X}^{2}$ is not monotone on the domain $\mathcal{S}_{d}^{+}$ and the matrix 4th power $\mathbf{X}\mapsto\mathbf{X}^{4}$ is neither monotone nor convex. The following are some of the few well-known nontrivial “operator-monotone” and “operator-convex” functions in the matrix analysis literature (cf. Bhatia [2013, Chapter V]).
Fact 2 (Operator-monotone and -concave Functions).
@@
[…]
Fact 3 (Trace Monotonicity).
@@
Fact 4 (Trace Jensen).
@@
2.2 Concentration of Scalars
We briefly review some classical and recent scalar tail bounds in this subsection. Markov inequality, the foundation of a large family of scalar concentration inequalities, states that if $X$ is a nonnegative random variable,
@@ 2.2 Concentration of Scalars
for any $a>0$. There have been many generalizations of (MI). Ville’s inequality [Ville, 1939] states that a similar tail bound holds uniformly over time for a nonnegative supermartingale $\{M_{n}\}$,
$\mathbb{P}(\exists n,\ M_{n}\geqslant a)\leqslant(\mathbb{E}M_{0})a^{-1},$ (VI)
or equivalently, $\mathbb{P}(M_{\tau}\geqslant a)\leqslant(\mathbb{E}M_{0})a^{-1}$ for any stopping time $\tau$ [Howard et al., 2020]. Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities (“confidence sequences”) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b] that have been favored by, for example, the multi-armed bandit literature.
Applying (MI) to the square of the sum of i.i.d. random variables $X_{1},\dots,X_{n}$ of mean $\mu$ and variance $\sigma^{2}$, we have the famous Chebyshev inequality
$\mathbb{P}\left(\left|\overline{X}_{n}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{na^{2}},$ (CI)
where $\overline{X}_{n}$ is the arithmetic mean of $X_{1},\dots,X_{n}$. A perhaps less-known extension of (CI) is that if $X_{1},\dots,X_{n}$ have central $p$th moment $v_{p}$ ($1\leqslant p\leqslant 2$) instead of variance,
$\mathbb{P}\left(\left|\overline{X}_{n}-\mu\right|\geqslant a\right)\leqslant\frac{2^{2-p}v_{p}}{n^{p-1}a^{p}},$ (5)
which we prove in Section A.1 for completeness. In the other direction, if $X_{i}$’s are light-tailed and have moment generating function (MGF) $G(\gamma)=\mathbb{E}\mathrm{e}^{\gamma(X_{1}-\mu)}$, (MI) applied to the exponential of their mean yields the Chernoff-Hoeffding inequality
$\mathbb{P}(\overline{X}_{n}-\mu\geqslant a)\leqslant G(\gamma/n)^{n}\mathrm{e}^{-\gamma a}.$ (6)
Recently, Ramdas and Manole [2023] discovered some interesting strengthened forms of the bounds above. First, uniform randomization. It is found that if $U$ is a $\mathrm{Unif}_{(0,1)}$-distributed random variable independent from $X$ or $\{M_{n}\}$, (MI) and (VI) and be losslessly tightened into
$\mathbb{P}(X\geqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (UMI)
and
$\mathbb{P}(M_{\tau}\geqslant Ua)\leqslant(\mathbb{E}M_{0})a^{-1}$ (UVI)
respectively. The downstream bounds including (CI), (5) and (6) can thus also be randomized into strictly tighter forms. Second, Ramdas and Manole [2023] introduce exchangeable generalizations of (CI) which partly shares the anytime virtue of (VI)-type bounds, stating that as long as the infinite sequence $X_{1},X_{2},\dots$ is exchangeable,
$\mathbb{P}\left(\exists n,\ \left|\overline{X}_{m}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{a^{2}};$ (XCI)
and further, if i.i.d., (XCI) implies that for any fixed $N$,
$\mathbb{P}\left(\exists n\geqslant N,\ \left|\overline{X}_{m}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{Na^{2}}.$ (7)
Extensions of (MI) and (VI) to random symmetric matrices are proved by Ahlswede and Winter [2002, Appendix C]; of (6) by Tropp [2012, 2015]. We will bring up these previous extensions again soon in the following. Our current paper is partly animated by the question if these novel bounds (UMI), (UVI), (XCI), and (7) by Ramdas and Manole [2023] have matrix siblings as well; and if yes, to what extent of generality. We shall show that there is indeed a plethora of matrix tail bounds that recover as $d=1$ special cases the results mentioned above, using classical matrix analysis tools plus some new probability techniques.
2.3 Templates of Our Bounds
We now sketch out the skeletons that our bounds may have, and discuss their relation. Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random symmetric matrices whose common distribution belongs to a family $\mathcal{P}$. Denote by $\mathbf{M}$ their common mean and $\overline{\mathbf{X}}_{n}$ their empirical average $\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. The following are the concentration events that we may prove to be of high probability. $\mathbf{A}\in\mathcal{S}^{+}_{d}$ denotes a constant matrix of deviation, and $r(\cdot)$ is a “randomization” map producing an $r(\mathbf{A})$ that is random but always smaller than $\mathbf{A}$ in Loewner order (the $Ua$ term in (UMI) is a one-dimensional example of $r(a)$). An arrow $A\to B$ denote the inclusion of events $A\subseteq B$, and hence also that $A$ is a stronger claim than $B$.
{tikzcd}
Most results in the past, for example (MHI), have been spectral bounds, which are weaker than many of the bounds we shall present in this paper.
3 Randomized Fixed-Time Matrix Tail Bounds
3.1 Randomized Matrix Markov Inequality
We begin our exposition with a matrix version of uniformly randomized Markov inequality (UMI) by Ramdas and Manole [2023]; a randomized version of the matrix Markov inequality by Ahlswede and Winter [2002, Theorem 12],

### v4

# Positive Semidefinite Matrix Supermartingales

@@ Abstract
We explore the asymptotic convergence and nonasymptotic maximal inequalities of supermartingales and backward submartingales in the space of positive semidefinite matrices. These are natural matrix analogs of scalar nonnegative supermartingales and backward nonnegative submartingales, whose convergence and maximal inequalities are the theoretical foundations for a wide and ever-growing body of results in statistics, econometrics, and theoretical computer science.
Our results lead to new concentration inequalities for either martingale dependent or exchangeable random symmetric matrices under a variety of tail conditions, encompassing now-standard Chernoff bounds to self-normalized heavy-tailed settings. Further, these inequalities are usually expressed in the Loewner order, are sometimes valid simultaneously for all sample sizes or at an arbitrary data-dependent stopping time, and can often be tightened via an external randomization factor.
@@ 1 Introduction
Let $\mathbf{Y}_{0},\mathbf{Y}_{1},\dots$ be a sequence of random symmetric $d\times d$ matrices adapted to a filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$. The process $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ is called a (matrix-valued) supermartingale if
for $\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}.$ (1)
Here, crucially, the order $\preceq$ is the Loewner order, where $\mathbf{A}\preceq\mathbf{B}$ if and only if $\mathbf{B}-\mathbf{A}$ is a positive semidefinite matrix. This introduces the rich spectral theory of matrices to the familiar $d=1$ scalar special case as well as the matrix martingale special case where the $\preceq$” in (1) takes the entry-wise “$=$”.
Similarly, if the matrices $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ are adapted instead to a backward filtration $\{\mathcal{G}_{n}\}_{n\geqslant 0}$, that is, $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$, it is called a (matrix-valued) backward submartingale if
for $\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n-1}|\mathcal{G}_{n})\succeq\mathbf{Y}_{n}.$ (2)
In this paper, we primarily study matrix-valued supermartingales and backward submartingales that are almost surely positive semidefinite (PSD). These matrix-valued stochastic processes are natural and non-trivial generalizations of scalar nonnegative supermartingales and nonnegative backward submartingales, studied extensively by the probability and statistics literature. In particular, as we shall review soon in Section 2.2, a scalar nonnegative supermartingale or backward submartingale $\{Y_{n}\}$ satisfies Ville’s inequality
$\mathbb{P}(\exists n,\ Y_{n}\geqslant a)\leqslant(\mathbb{E}Y_{0})a^{-1},$ (3)
which is interconnected with various results on the asymptotic and non-asymptotic behavior of these classes of processes, and has led to numerous methods in sequential statistics. We extend all these results and applications from scalars to matrices. In particular, we present the matrix Ville’s inequality
$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1})$ (4)
for both classes of processes, their numerous equivalent or strengthened forms, and a variety of statistical applications.
A list of major contributions of this paper is as follows.
@@
    The convergence and optional stopping theorems, and the Ville’s inequality for PSD matrix supermartingales (Theorems 3.2, 3.3 and 3.4) that generalize the corresponding scalar results.
@@
    Sequential testing procedure with a matrix test supermartingale (Algorithm 1) that can be applied to the meaning testing of bounded (Examples 3.7 and 3.10), symmetric (Example 3.9), or heavy-tailed (Examples 3.8 and 3.12) martingale-dependent matrices.
@@
    Matrix e-values and e-processes (Definitions 3.13 and 3.14), generalizing the corresponding scalar concepts that have recently been extensively studied.
@@
    The matrix Ville’s inequality for backward PSD matrix submartingales (Theorem 4.3), leading to the exchangeable matrix Chebyshev’s inequality (Theorem 4.5).
@@
    The randomized matrix Markov’s and Ville’s inequalities (Theorems 5.3 and 5.5).
  6. 6.
    A sequential mean testing example with the matrix supermartingale outperforming the scalar supermartingale (Example 6.5).
The rest of the paper is organized as follows. We review in Section 2 the technical background of matrix analysis, sequential statistics with scalar processes in the abovementioned classes, as well as matrix concentration inequalities in the literature. We study forward PSD matrix supermartingales, their Ville’s inequality and applications in Section 3. We then study backward PSD matrix submartingales, their Ville’s inequality and applications in Section 4. In Section 5 we provide several perspectives comparing our inherent matrix approach to some existing scalar-born or scalarized results, and discuss a randomized improvement that is further developed in Appendix A. An interesting empirical study is presented in Section 6, including a problem where our matrix method yields higher out-of-the-box power compared to the previous scalar method. Omitted proofs are found in Appendix B; some further references and independent theoretical results in Appendix C.
2 Preliminaries and Related Work
@@ 2.1 Matrix Analysis
Let $\mathcal{S}_{d}$ denote the set of all $d\times d$ real-valued symmetric matrices, which is the only class of matrices considered in this paper. [^1] We only briefly discuss the case with non-symmetric and non-square matrices in Section C.3. These matrices are denoted by bold upper-case letters $\mathbf{A},\mathbf{B}$, etc; whereas $d$ dimensional (column) vectors by bold lower-case letters $\mathbf{u},\mathbf{v}$, etc. For $I\subseteq\mathbb{R}$, we denote by $\mathcal{S}_{d}^{I}$ the set of all real symmetric matrices whose eigenvalues are all in $I$. $\mathcal{S}_{d}^{[0,\infty)}$, the set of positive semidefinite and $\mathcal{S}_{d}^{(0,\infty)}$, the set of positive definite matrices are simply denoted by $\mathcal{S}_{d}^{+}$ and $\mathcal{S}_{d}^{++}$ respectively. We consider random matrices taking values in $\mathcal{S}_{d}$. It is easy to see that if $\mathbf{X}$ is an integrable random matrix taking values in $\mathcal{S}_{d}^{I}$ where $I$ is an interval, any conditional expectation $\mathbb{E}(\mathbf{X}|\mathcal{A})$ also takes values in $\mathcal{S}_{d}^{I}$. The Loewner partial order $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite, and $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite.
Any function $f:I\to J$ (where $I,J\subseteq\mathbb{R}$) is identified canonically with a function $f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$ in the following way: for a diagonal matrix $\mathbf{\Sigma}\in\mathcal{S}_{d}^{I}$, $f(\mathbf{\Sigma})$ is obtained by applying $f$ entrywise; otherwise it is obtained by applying $f$ after orthogonal diagonalization, $f(\mathbf{U}^{\mathsf{T}}\mathbf{\Sigma}\mathbf{U})=\mathbf{U}^{\mathsf{T}}f(\mathbf{\Sigma})\mathbf{U}$. Notable matrix-to-matrix operators that recur in this paper defined this way include the matrix exponential $\exp:\mathcal{S}_{d}\to\mathcal{S}^{++}_{d}$, the matrix logarithm $\log:\mathcal{S}^{++}_{d}\to\mathcal{S}_{d}$, the matrix absolute value [^2] We use $\operatorname{abs}(\,\cdot\,)$ instead of ||$\cdot$|| ⋅ | to distinguish it from the determinant and other norms. $\operatorname{abs}:\mathcal{S}_{d}\to\mathcal{S}^{+}_{d}$, and powers $\mathbf{X}\mapsto\mathbf{X}^{k}$ for any $k\in\mathbb{R}$.
@@
It is rare to find matrix-to-matrix functions that are monotone (i.e. preserving the $\preceq$ order) and convex (i.e. $f$ such that $f(\sum\lambda_{i}\mathbf{X}_{i})\preceq\sum{\lambda_{i}}f(\mathbf{X}_{i})$ if $\sum\lambda_{i}=1$ and each $\lambda_{i}\geqslant 0$; concavity is defined similarly). Scalar-to-scalar monotonicity ($I\to J$) does not imply matrix-to-matrix monotonicity ($\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$), and the same with convexity. For example, the matrix square $\mathbf{X}\mapsto\mathbf{X}^{2}$ is not monotone on the domain $\mathcal{S}_{d}^{+}$ and the matrix 4th power $\mathbf{X}\mapsto\mathbf{X}^{4}$ is neither monotone nor convex. The following are some of the few well-known nontrivial “operator monotone” and “operator convex” functions in the matrix analysis literature (cf. Bhatia [2013, Chapter V]).
Fact 1 (Operator Monotone and Convex Functions).
@@
With trace, however, real-to-real functions translate into matrix-to-real functions with monotonicity and convexity intact. That is,
Fact 2 (Trace Monotonicity).
@@
Fact 3 (Trace Jensen).
@@
2.2 Scalar Nonnegative Supermartingales and Statistical Applications
A scalar-valued nonnegative supermartingale $\{Y_{n}\}_{n\geqslant 0}$ on the filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ satisfies Ville’s inequality [Ville, 1939],
$\mathbb{P}(\exists n,\ Y_{n}\geqslant a)\leqslant(\mathbb{E}Y_{0})a^{-1},$ (VI)
or equivalently,
$\mathbb{P}(Y_{\tau}\geqslant a)\leqslant(\mathbb{E}Y_{0})a^{-1}\quad\text{for any stopping time }\tau.$ (VI)
The equivalence is demonstrated by, for example, Howard et al. [2021, Lemma 3]. Ville’s inequality is often seen as a time-uniform, or a “stopped” generalization of Markov’s inequality
@@ 2.2 Scalar Nonnegative Supermartingales and Statistical Applications
for a nonnegative random variable $X$.
As Markov’s inequality lays the foundation for a large body of concentration inequalities for a fixed number of random variables leading up to non-asymptotic statistical methods (e.g. confidence intervals) valid at fixed finite sample sizes, Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities and accordingly sequential statistical methods which are valid at data-dependent sample sizes. For example, a $(1-\alpha)$-confidence sequence, which is derived almost exclusively via (VI) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b], is a sequence of confidence sets $\{C_{n}\}_{n\geqslant 1}$ for the parameter $\theta$ such that $\mathbb{P}(\theta\in C_{n}\text{ for all }n)\geqslant 1-\alpha$, or equivalently $\mathbb{P}(\theta\in C_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$ that can possibly depend on the data. More generally, the stopped version of (VI) is used in sequential hypothesis testing, as it implies that to construct a conservative sequential test of $\alpha$ level it suffices to construct a process $\{Y_{n}\}$ with $Y_{0}=1$ that is a nonnegative supermartingale under the null hypothesis, and reject it when $Y_{n}$ first surpasses $1/\alpha$. Various examples of sequential tests constructed via nonnegative supermartingales can be found in the recent review article by Ramdas et al. [2023, Sections 4 and 5]. We return to this topic of test processes with a more involved discussion in Section 3.4.
Numerous variations and extensions on (VI) exist. First, the backward Ville’s inequality (see e.g. Manole and Ramdas [2023, Theorem 2]). A process $\{Y_{n}\}$ adapted to a backward filtration $\{\mathcal{G}_{n}\}$ is called a backward submartingale if $\mathbb{E}(Y_{n-1}|\mathcal{G}_{n})\geqslant Y_{n}$ which, if nonnegative, satisfies (VI) as well,
$\mathbb{P}(\exists n,\ Y_{n}\geqslant a)\leqslant(\mathbb{E}Y_{0})a^{-1}.$ (7)
Second, the uniformly randomized Ville’s inequality [Ramdas and Manole, 2024, Theorem 4.1], which states that (VI) can be tightened into
$\mathbb{P}(Y_{\tau}\geqslant Ua)\leqslant(\mathbb{E}Y_{0})a^{-1}\quad\text{forany stopping time }\tau,$ (8)
where $U\sim\operatorname{Unif}_{(0,1)}$ is independent from the filtration $\{\mathcal{F}_{n}\}$. Ramdas and Manole [2024, Section 4.3] outline its implication for sequential statistics, which we discuss in Section 5.3 in the context of matrices.
2.3 Matrix Concentration Inequalities
2.3.1 Spectral Bounds
There has been a wealth of results bounding the tails of the largest eigenvalue (or the spectral norm) of the sum of independent or martingale dependent random symmetric matrices. Most notably, Oliveira [2009, 2010], Tropp [2011, 2012, 2015], and their collaborators achieve the technically involved extension of the Cramér-Chernoff method, the central tool in deriving exponentially decaying scalar concentration inequalities, to random matrices satisfying similar tail conditions. The most prominent step along this route is the application of Lieb’s concavity theorem [Lieb, 1973] that ends up bounding the expected value of the scalar
$\mathrm{tr}\exp\left(\sum\mathbf{X}_{i}\right)$ (9)
where $\mathbf{X}_{i}$’s are independent [Tropp, 2012, Section 3.4] or martingale dependent [Tropp, 2011, Section 2.2] matrices. Fixed-time tail bounds follow from Markov’s inequality (MI). Already noticed by Tropp [2011] and further exploited by Howard et al. [2020], the trace process (9) is a (scalar) nonnegative supermartingale, and Ville’s inequality (VI) thus implies time-uniform spectral tail bounds for the same classes of random matrices.
Our new results will take a different approach as we shall construct matrix-valued supermartingales and Loewner-ordered ($\preceq$) concentration inequalities. Their fixed-time counterparts are reviewed next in Section 2.3.2. Our results can often recover these “scalarized” results of Oliveira, Tropp, and Howard et al. mentioned above.
2.3.2 Ahlswede-Winter Bounds
Ahlswede and Winter [2002, Theorem 12] prove the following Loewner-ordered matrix Markov’s inequality